OpenStudy (anonymous):

Simplify (x2 + 5x + 12)(x2 + 6x + 7). A. x4 + 37x2 + 84 B. x4 + 30x2 + 84 C. x4+ 11x3 + 49x2 + 107x + 84 D. x4 + 6x3 + 7x2 + 11x + 19 is it C?

OpenStudy (anonymous):

(x(2)+5x+12)(x(2)+6x+7) =(x(2)+5x+12)(x(2)+6x+7) =(x(2))(x(2))+(x(2))(6x)+(x(2))(7)+(5x)(x(2))+(5x)(6x)+(5x)(7)+(12)(x(2))+(12)(6x)+(12)(7) =4x2+12x2+14x+10x2+30x2+35x+24x+72x+84 =56x2+145x+84 I think you are right hon.

OpenStudy (anonymous):

:D thank you!

OpenStudy (anonymous):

i need help with one last problem will you help me?